OpenStudy (anonymous):

verify (cosA-cosB)/(sinA+sinB) = -tan [(A-B)/2]

OpenStudy (anonymous):

\[\cos C-\cos D=-2 \sin \frac{ C+D }{ 2 }\sin \frac{ C-D }{ 2 }\] \[\sin C+\sin D=2 \sin \frac{ C+D }{ 2 }\sin \frac{ C-D }{ 2 }\]